Circular No. 8856

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

CBAT@CFA.HARVARD.EDU (science)

URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 N2 = P/2002 O5 (NEAT)

Comet P/2002 O5 (cf. IAUC 7942) has been recovered apparently independently on CCD images obtained at three different observatories:

2007 UT α_{2000} δ_{2000} Mag. Observer
July 15.22509 17^{h}05^{m}36.^{s}69 −12^{o}42′18′′.5 15.8 LINEAR
15.23619 17 05 38.35 −12 41 00.0 16.0  _{”}
15.24728 17 05 39.89 −12 39 40.4 16.6  _{”}
15.25845 17 05 41.73 −12 38 23.8 15.9  _{”}
15.26954 17 05 43.38 −12 37 05.8 15.9  _{”}
16.00235 17 07 43.80 −11 11 33.0 16.4 Fratev
16.00372 17 07 44.31 −11 11 26.6 17.0 Lombardi
16.00635 17 07 44.71 −11 11 07.8 17.1  _{”}
16.00872 17 07 45.02 −11 10 49.7 16.5 Fratev
16.01498 17 07 45.87 −11 10 06.3 16.3  _{”}
16.02155 17 07 46.73 −11 09 19.2 16.5  _{”}

LINEAR (Socorro, NM). 1.0-m f/2.15 reflector. Observations identified via

the normal checking of data by the Minor Planet Center. G. Lombardi and E. Pettarin (Farra d’Isonzo, Italy). 0.40-m f/4.5 reflector.

USNO-A2.0 catalogue. Comet has a diffuse coma of diameter ≈ 10′′ and a tail ≈ 15′′ long in p.a. 135^{o}.

F. Fratev and E. Mihaylova (Plana, Bulgaria). 0.25-m f/3.2 reflector. USNO-B1.0 catalogue. Coma diameter ≈ 25′′; total red mag 16.0.

The correction to the predictions on MPC 51823 and in the 2007 Comet Handbook is ΔT = −0.36 day. The following linked orbital elements are by B. G. Marsden, from 254 observations, 2002–2007 (mean residual 0′′.7):

Epoch = 2002 July 25.0 TT

T = 2002 Aug. 3.10144 TT  ω = 15.^{o}30571
e = 0.5973957  Ω = 282.20995 }2000.0
q = 1.1742394 AU  i = 20.39966

a = 2.9166089 AU n^{o} = 0.19787311 P = 4.98 years

Epoch = 2007 Aug. 8.0 TT

T = 2007 July 25.90695 TT  ω = 15.^{o}27124
e = 0.5976949  Ω = 282.20178 }2000.0
q = 1.1733898 AU  i = 20.40165

a = 2.9166666 AU n^{o} = 0.19786724 P = 4.98 years

2007 July 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green